OpenStudy (anonymous):

Find the axis of symmetry of the parabola formed by the following quadratic function: y =-5x^2+7x-9 A. x = -7/5 B. x = 7/5 C. x = -7/10 D. x = 7/10

OpenStudy (cwrw238):

first divide through by -5 then convert to vertex form

OpenStudy (anonymous):

how do i do that ? @cwrw238

OpenStudy (cwrw238):

divide by -5 ;- =-5(x^3 -7x/5 + 9/5) = -5 ((x - 7/10)^2 +9/5 - 49/100)) x = 7/10 is the axis of symmetry

OpenStudy (cwrw238):

another way to do this is to use the formula x = -b/2a from the general form a^x^2 + bx + c = 0 b = 7 and 2a = 2*-5 = -10 x = -7/-10 x = 7/10